# MeshArt: Generating Articulated Meshes with Structure-guided Transformers

### v1

@@ 1 Introduction
To address these challenges, we introduce MeshArt111daoyig.github.io/Mesh_Art/, a hierarchical transformer-based model that generates articulated 3D meshes with sharp details and an efficient triangle mesh representation, resembling the compactness of artist-created 3D models. Our key insight is to decompose articulated meshes into high-level object structures and low-level part geometries: high-level structures are represented by part bounding primitives endowed with articulation properties, while part geometries are generated based on both articulation structures and local mesh connectivity. To enable effective training of MeshArt, we enhance PartNet [21], the largest 3D object dataset with part labels, by adding articulation joint annotations to the table, chair, and storage categories, increasing the number of articulated objects 6×6$\times$6 × compared to existing articulation datasets [6, 38].
@@ 2 Related Work
Mesh Generation. Several methods have been proposed to directly generate mesh representation, due to their compact nature more similar to human-created 3D models. Earlier approaches such as Scan2Mesh [3] proposed to generate mesh structures as graphs, but were limited to generation of small meshes with very limited number of vertices. PolyGen [22] proposed to employ two separate auto-regressive transformers to learn the distributions of vertices and faces independently. Similarly, Polydiff [1] uses a diffusion model to generate triangle soups to construct meshes. MeshGPT [28] instead tokenizes mesh representations with a graph neural network (GNN)-based encoder and learns mesh tokens using a GPT-style transformer. PivotMesh [33] enhances such generation by guiding the process with pivot vertices that define the coarse shape of an object.
Despite these advancements, the functionality of the generated assets is often overlooked, as these methods all focus on static mesh generation. In contrast, our approach disentangles structural part semantics from geometry, employing an autoregressive model to facilitate the generation of functional articulated objects.
Structured Object Generation. Structure-aware generative models for 3D objects have focused on modeling object parts and their relationships. SAGNet [36] uses a VAE to encode part geometry and pairwise relations into a latent space, which is then decoded to generate coherent objects. GRASS [13] models part structure hierarchies as binary trees based on symmetry hierarchies [32], utilizing a recursive autoencoder to capture both parts and their geometric properties. StructureNet [20] extends this idea by representing objects as n-ary graphs, enabling the generation of more complex structures.
SDM-NET [5] combines deformable mesh boxes to represent part geometry and a VAE to encode the global structure, producing more detailed part geometries. DSG-Net [39] disentangles part geometry from structure by learning separate latent spaces for each, allowing fine-grained control during generation.
@@
While these methods effectively generate coherent 3D shapes with structural parts, they do not account for articulation. In contrast, we encode articulation directly within the structural generation of triangle meshes, enabling the creation of compact, functional meshes with articulated properties.
@@
Articulated Object Modeling. Recent advancements in articulated 3D object modeling have focused on reconstructing or generating objects with geometric and motion properties. Ditto [10] reconstructs articulated objects from point clouds as implicit representations, capturing geometric occupancy, part types, and joint parameters. Similarly,  [9] and [16] reconstruct articulated objects as implicit neural fields from stereo RGB and point cloud videos, respectively. PARIS [14] disentangles static and movable parts from two observations at different articulation states, predicting both geometry and motion parameters. Building on this, [34] introduces explicit point-level correspondence between reconstructed shapes in two states, enabling the modeling of more complex articulated objects.
@@
NAP [12] introduces the unconditional generation of articulated objects by modeling articulation parameters with geometry using an articulation tree. The final geometry is decoded either via an implicit field decoder or through part retrieval. CAGE [15] employs diffusion to generate articulation abstractions, relying on part retrieval for final object geometry. However, this part retrieval and assembly approach [41, 35, 30] can introduce inconsistencies in part geometry and overall shape structure. In contrast, our MeshArt generates both structure and geometry cohesively in a sharp, unified triangle mesh representation, enhancing the efficiency and robustness of articulated object modeling.
@@ 3.2 Structure-guided Part Mesh Generation
The input sequence for $\Phi_{g}$ is constructed by placing the structure and junction tokens at the beginning, followed by the triangle tokens $\mathcal{T}_{g}^{i}$. Concretely, for the current part mesh $\mathcal{P}^{i}$, its corresponding structure tokens $(t^{i}_{1},\ldots,t^{i}_{12})\subset\mathcal{T}_{s}$ are extracted and cross-attended to the overall object structure $\mathcal{T}_{s}$, before being added to the start of the sequence. Next, we retrieve ground-truth junction faces from other parts and append their tokens to the sequence, prior to triangle tokens $\mathcal{T}_{g}^{i}$. These junction tokens inform smooth transitions and connectivity between neighboring parts.
@@ 3.3 Hierarchical Articulated Mesh Sampling
For articulated mesh generation at inference time, we first sample the structure transformer $\Phi_{s}$ unconditionally to generate an object structure sequence $\mathcal{S}$ in an autoregressive manner. The structure decoder $\mathbf{D}_{s}$ then decodes $\mathcal{S}$ into part bounding boxes with associated articulation properties.
Next, for each part $\mathcal{P}^{i}$, we generate its mesh triangle sequence $\mathcal{T}_{g}^{i}$ by sampling the geometry transformer $\Phi_{g}$, conditioned on object structure $\mathcal{S}$ and predicted junction faces from previously generated part meshes. Only junction faces that are spatially near the bounding box $\mathbf{b}^{i}$ of $\mathcal{P}^{i}$ are considered. The predicted triangle tokens $\mathcal{T}_{g}^{i}$ are then decoded into mesh triangles by the geometry decoder $\mathbf{D}_{g}$. Finally, the object mesh is obtained by merging close vertices of the generated triangles. This articulated mesh can be transformed into different articulation states based on the structure-level joint predictions.
@@ 5 Experiments
Shape Quality. Evaluating the unconditional synthesis of articulated 3D shapes presents challenges due to the lack of direct ground-truth correspondence; several evaluation metrics have been proposed to compare the distribution of generated and ground-truth objects. We follow prior work [12, 15], and evaluate Instantiation Distance (ID) and Abstract Instantiation Distance (AID). ID evaluates the articulated mesh geometry at different articulation states, while AID evaluates only the object structure at these states. Following [15], we synchronize the articulation states of both generated and ground-truth meshes, using 10 evenly spaced articulation states between the start and end positions. We sample 2,048 points on the surface of the generated mesh or its structure bounding box for ID and AID, respectively. We then compute Minimum Matching Distance (MMD), Coverage (COV), and 1-Nearest-Neighbor Accuracy (1-NNA). Lower MMD indicates better mesh quality, higher COV reflects broader coverage of the real data distribution; for 1-NNA, 50% is optimal.
Perceptual Quality. The aforementioned metrics measure geometric shape quality, but do not address the visual similarity of the generated meshes to the real distribution. To assess this, we render both generated and ground-truth meshes at the same articulation state from eight different viewpoints using Blender, applying a metallic material to emphasize detailed geometric structures. Subsequently, we calculate FID (Fréchet Inception Distance) and KID (Kernel Inception Distance) scores for these image sets. For both FID and KID, lower scores indicate better performance.
@@ 5.2 Results
Figure 4: Qualitative comparison across all three categories in our dataset, showing the generated structure with joints, the mesh geometry in its resting state, and an articulated state. Our approach produces compact meshes with sharp geometry, and precise joint generation enables more realistic part movements.
@@
Shape Novelty Analysis. We conduct an analysis of our generated shape novelty following [33, 28]. A total of 1000 shapes are generated, and their nearest neighbors in the training set are retrieved based on Chamfer Distance (CD). As shown in Fig. 5, the nearest neighbor analysis demonstrates that our method not only well-captures the training shape distribution (evidenced by low CD values), but moreover, effectively generates novel and realistic shapes (indicated by higher CD values).
Figure 5: Shape novelty analysis on storage category. We generate 1000 shapes with our method and plot the distribution of their distance to the training set. We visualize the generated shape (left) and its closest train shape (right) at different distance percentiles. The generated shape at ${}^{\text{th}}$ percentile already differs from the closest train shape. This novelty analysis shows that our method can both cover shape distribution (low CD) and generate novel shapes (high CD).
@@ 6 Conclusion
We introduced MeshArt, a triangle-based, cohesive hierarchical approach for generating articulated 3D meshes. We decompose objects into coarse part-based structures and detailed part mesh geometries, grounding articulations at the structure level. This enables synthesizing clean, coherent articulated meshes, ensuring both geometric quality and functional properties. We believe MeshArt will drive progress in 3D vision and embodied AI, opening new possibilities for digital agents to learn and interact with the physical world.
@@
In this supplementary document, we provide additional details about MeshArt. In Sec. 7, we give more implementation details of our method and loss functions. We elaborate our data annotation process in Sec. 8. We also include additional quantitative comparisons in Sec. 9. We highly encourage readers to watch the supplemental video to see more articulated object generations in action.
@@ 7.1 Structure VQ-VAE
The decoder will output a set of joint information per-triangle. To obtain a single set of joint predictions per part, the outputs are averaged across all triangles within the part.
This architecture effectively learns quantized embeddings for articulated object structures. These embeddings serve as the basis for the structure transformer, enabling autoregressive generation of object structures with articulations.
@@ 9 Additional Results
We compare the mesh generation quality of our method with NAP [12], and the state-of-the-art direct mesh generation approach, MeshGPT [28]. Since MeshGPT does not predict object part and articulation information, the evaluation is performed on generated meshes in their canonical resting state. As shown in Tab. 6, we also achieve comparable COV scores to MeshGPT while outperforming the MMD score, indicating higher fidelity in the generated shapes. Notably, our method shows significant mesh generation improvement over NAP on all metrics.

### v2

@@ 1 Introduction
To address these challenges, we introduce MeshArt [^1] Project page: daoyig.github.io/Mesh_Art/, a hierarchical transformer-based model that generates articulated 3D meshes with sharp details and an efficient triangle mesh representation, resembling the compactness of artist-created 3D models. Our key insight is to decompose articulated meshes into high-level object structures and low-level part geometries: high-level structures are represented by part bounding primitives endowed with articulation properties, while part geometries are generated based on both articulation structures and local mesh connectivity. To enable effective training of MeshArt, we enhance PartNet [24], the largest 3D object dataset with part labels, by adding articulation joint annotations to the table, chair, and storage categories, increasing the number of articulated objects 6×6$\times$6 × compared to existing articulation datasets [8, 43].
@@ 2 Related Work
Mesh Generation. Several methods have been proposed to directly generate mesh representation, due to their compact nature more similar to human-created 3D models. Earlier approaches such as Scan2Mesh [5] proposed to generate mesh structures as graphs, but were limited to small meshes with few vertices. PolyGen [25] used two autoregressive transformers to model vertex and face distributions separately, while Polydiff [1] applied diffusion to generate triangle soups. MeshGPT [32] tokenized meshes via a GNN-based encoder and learned mesh token generation with a GPT-style transformer. PivotMesh [38] improved generation by guiding it with pivot vertices to define coarse object shapes. More recent approaches, including MeshAnythingV2 [3] and EdgeRunner [35], tokenized meshes by forming face sequences from adjacent vertices. MeshXL [2] introduced spatial coordinate embeddings, while Meshtron [10] employed an hourglass transformer for high-resolution mesh generation.
In contrast to these advances in static mesh generation, our work focuses on generating functional articulated objects by disentangling structure from geometry with hierarchical autoregressive models.
Structured Object Generation. Structure-aware generative models for 3D objects have focused on modeling object parts and their relationships [16, 37]. SAGNet [41] learns part geometry and pairwise relations within a shared latent space, while StructureNet [23] represents objects as n-ary graphs, using a hierarchical network to encode both aspects. SDM-NET [7] models part geometry with deformable mesh boxes and encodes global structure in a VAE. DSG-Net [44] learns separate latent space for geometry and structure, achieving fine-grained control during generation.
@@
While these methods effectively generate coherent 3D shapes with structural relations, they do not model articulation. In contrast, we encode articulation properties at the object structure level, enabling the creation of compact meshes with functional part motions.
@@
Articulated Object Modeling. Recent advances in articulated 3D object modeling focus on reconstruction or generation with geometric and motion properties. Ditto [13] reconstructs articulated objects from point clouds as implicit representations, capturing occupancy, part types, and joint parameters. Similarly,  [12] and  [19] use implicit neural fields to reconstruct articulation from stereo RGB and point cloud videos. PARIS [17] disentangles static and movable parts from two articulation states, predicting both geometry and motion. Building on this, [39] introduces explicit point-level correspondence between articulated states, enabling more complex object modeling.
@@
NAP [15] models 3D objects with an articulation tree, decoding final geometry via implicit fields or part retrieval. CAGE [18] employs diffusion to generate articulation abstractions, relying on part retrieval for final output. However, retrieval-based approaches [47, 40, 34] can bring inconsistencies in part geometry and overall structure. In contrast, our approach generates structure and geometry cohesively in a unified triangle mesh representation, improving efficiency and robustness in articulated object modeling.
@@ 3.2 Structure-Guided Part Mesh Generation
The input sequence for $\Phi_{g}$ is constructed by placing the structure and junction tokens at the beginning, followed by the triangle tokens $\mathcal{T}_{g}^{i}$. Concretely, for the current part mesh $\mathcal{P}^{i}$, its corresponding structure tokens $(t^{i}_{1},\ldots,t^{i}_{12})\subset\mathcal{T}_{s}$ are extracted and cross-attended to the overall object structure $\mathcal{T}_{s}$, before being added to the start of the sequence. Next, the junction faces from other parts are appended to the sequence before triangle tokens $\mathcal{T}_{g}^{i}$, informing smooth transitions and connectivity between neighboring parts.
@@ 3.3 Hierarchical Articulated Mesh Sampling
At inference time, we first sample the structure transformer $\Phi_{s}$ unconditionally to generate an object structure sequence $\mathcal{S}$, which is then decoded by the structure decoder $\mathbf{D}_{s}$ into part bounding boxes with articulation properties.
Next, for each part $\mathcal{P}^{i}$, we generate its triangle sequence $\mathcal{T}_{g}^{i}$ by sampling the geometry transformer $\Phi_{g}$, conditioned on object structure $\mathcal{S}$ and junction faces from nearby part meshes. We cache all the triangles predicted as junctions by $\mathbf{D}_{g}$, then retrieve the junction faces from the cache based on their proximity to the current structure bounding box $\mathbf{b}^{i}$.
The predicted triangle tokens $\mathcal{T}_{g}^{i}$ are then decoded into mesh triangles by the geometry decoder $\mathbf{D}_{g}$. Finally, the object mesh is obtained by merging close vertices of the generated triangles. This articulated mesh can be transformed into different articulation states based on the joint predictions.
@@ 5 Experiments
Shape Quality. Following [15, 18], we use Instantiation Distance (ID) to assess mesh geometry at different articulation states and Abstract Instantiation Distance (AID) to evaluate object structure at these states. As in [18], we synchronize articulation states of both generated and ground-truth meshes by sampling 10 evenly spaced articulation states between start and end positions. To evaluate ID and AID, we sample 2,048 surface points from each object mesh and its structure bounding box, respectively. We then compute Minimum Matching Distance (MMD), Coverage (COV), and 1-Nearest-Neighbor Accuracy (1-NNA). Lower MMD indicates better quality, higher COV reflects broader distribution coverage; for 1-NNA, 50% is optimal.
Perceptual Quality. While the metrics above assess geometric quality, they do not capture visual similarity to the real data distribution. To address this, we render generated and ground-truth meshes at the same articulation state from 8 viewpoints in Blender, using a metallic material to highlight geometric details. We then compute FID (Fréchet Inception Distance) and KID (Kernel Inception Distance), where lower scores indicate better performance.
@@ 5.2 Results
Figure 4: Qualitative comparison across all categories, showing generated structures with joints, mesh geometry in resting and articulated states. Our approach generates compact meshes with sharp geometry and precise joints, enabling realistic part movements.
@@
Shape Novelty Analysis. We analyze the shape generation novelty following [38, 32]. We generate 1,000 shapes and retrieve their nearest neighbors in the training set using Chamfer Distance (CD). As shown in Fig. 5, our method captures the training distribution well (indicated by low CD) while effectively generating novel shapes (with high CD).
Figure 5: Shape novelty analysis on storage category. We generate 1,000 shapes and plot their distances to the train set. The generated shape (left) at ${}^{\text{th}}$ percentile already differs from its closest match (right), showing that our method captures both the shape distribution (low CD) while generating novel shapes (high CD).
Conditional Generation. Our approach can support conditional generation by injecting condition tokens to the structure transformer $\Phi_{s}$ while keeping the geometry transformer $\Phi_{g}$ fixed. As shown in Fig. 6, our approach generates articulated meshes from 3D point clouds and sketch images. More results can be found in the supplemental.
Figure 6: Conditional Generation. Articulated structures and geometry are generated conditioned on point clouds or sketches.
@@ 6 Conclusion
We introduced MeshArt, a triangle-based hierarchical approach for generating articulated 3D meshes. We decompose objects into coarse part-based structures and detailed part mesh geometries, grounding articulations at the structure level. This enables synthesizing clean, coherent articulated meshes, ensuring both geometric quality and functional properties. We believe MeshArt will drive progress in 3D vision and embodied AI, opening new possibilities for digital agents to learn and interact with the physical world.
@@
In this supplementary document, we provide additional details about MeshArt. In Sec. 7, we give more implementation details of our method and loss functions. We elaborate our data annotation process in Sec. 8. We also include additional quantitative comparisons in Sec. 9. We encourage readers to watch the supplemental video to see more articulated object generations in action.
@@ 7 Method Details
We use VQVAEs to model both part articulations and mesh geometries for our hierarchical transformers. Our structure VQVAE encodes extra part-level features (e.g., semantics, geometry feature, and articulation joint), alongside vertex locations, into a compact latent space for articulation-aware structure generation, while our geometry VQVAE predicts additional junction face probabilities for coherent part mesh generation.
@@ 7.1 Structure VQ-VAE
The decoder will output a set of joint information per triangle. To obtain a single set of joint predictions per part, the outputs are averaged across all triangles within the part.
This architecture effectively learns quantized embeddings for articulated object structures. These embeddings serve as the basis for the structure transformer, enabling the autoregressive generation of object structures with articulations.
@@ 8 Data Annotation
Figure 9: Our method can generate sharp geometry and realistic articulations for microwaves.
Figure 10: Conditional Generation. Articulated structures and geometry are generated conditioned on point clouds or sketches.
9 Additional Results
Quantitative Comparison at Resting State. We compare the mesh generation quality of our method with NAP [15], and the state-of-the-art direct mesh generation approach, MeshGPT [32]. Since MeshGPT does not predict object part and articulation information, the evaluation is performed on generated meshes in their canonical resting state. As shown in Tab. 6, we also achieve comparable COV scores to MeshGPT while outperforming the MMD score, indicating higher fidelity in the generated shapes. Notably, our method shows significant mesh generation improvement over NAP on all metrics.
Additional Categories. We show additional results on microwaves in Fig. 9.
Qualitative Comparison with CAGE. We compare MeshArt with CAGE [18], which retrieves part geometry based on conditionally generated articulated structures. As shown in Fig. 11, our method achieves coherent shape synthesis while achieving sharp geometry details, avoiding undesired part collision and inconsistencies between parts.
Figure 11: Visual comparison with CAGE. Our method can generate coherent shapes with sharp geometry.
Conditional Generation. MeshArt can generate articulated structure and geometry conditioned on point clouds or sketch images. We extract input features using Michelangelo [45] for 3D point clouds and Radio [29] for sketches. A linear layer projects these features to match the structure transformer $\Phi_{s}$’s feature space, appending them to the beginning of the structure sequence. The transformer $\Phi_{s}$ then learns to generate articulation-aware structure bounding boxes. Since the structure codebook $\mathcal{C}_{s}$ remains fixed, the geometry transformer $\Phi_{g}$ requires no finetuning. As shown in Fig. 10, given the 3D/2D conditions, our method can generate plausible articulated structures, which then guide faithful synthesis of part geometry.